# Exploring the Interplay between Musical Preferences and Personality through the Lens of Language

### v1

@@ Abstract
Music serves as a powerful reflection of individual identity, often aligning with deeper psychological traits. Prior research has established correlations between musical preferences and personality traits, while separate studies have demonstrated that personality is detectable through linguistic analysis. Our study bridges these two research domains by investigating whether individuals’ musical preferences are recognizable in their spontaneous language through the lens of the Big Five personality traits (Openness, Conscientiousness, Extroversion, Agreeableness, and Neuroticism). Using a carefully curated dataset of over 500,000 text samples from nearly 5,000 authors with reliably identified musical preferences, we build advanced models to assess personality characteristics. Our results reveal significant personality differences across fans of five musical genres. We release resources for future research at the intersection of computational linguistics, music psychology and personality analysis.
@@ 1 Introduction
"Our language is the reflection of ourselves.
A language is an exact reflection of the character and growth of its speakers."
Cesar Chavez
Music is a powerful expression of individuality, often reflecting deeper aspects of one’s personality. The relationship between musical preferences and personality traits has long been a subject of interest in psychology [7, 29, 6, 11]. Prior studies indicate that certain genres are associated with specific personality traits, suggesting that music genre preferences offer valuable insights into individuals’ psychological profile [6, 23, 4].
@@
Our study aims to connect the two lines of research – we explore the relationship between musical preferences and personality traits, as manifested in authentic written language. Specifically, we ask if there are detectable (similarities and) differences in the Big Five personality traits in the language of people with various musical preferences, when authoring spontaneous textual content on social media. We consider five popular and diverse musical genres: Classical music, Hip-Hop, Metal, Indie and Electronic. The hypothesis driving this research posits that despite the inherent complexity of automatic personality analysis, the differences in personality traits of people with various musical preferences, "shine through" their language, to the extent that can be automatically captured through advanced natural language processing techniques.
We explore the hypothesis on a carefully curated large and diverse dataset, consisting of over 500,000 text samples written by nearly 5,000 authors with reliably identified musical preferences. These samples were collected from a range of public non-music-related online forums, ensuring a broad topical representation. Each participant was linked to a single dominant music genre they frequently engage with, enabling unbiased textual personality extraction and a focused analysis of the relationship between personality traits and musical preferences. Leveraging text generation capabilities of contemporary LLMs we then gather high-quality data for training personality detectors along the five dimensions: Openness, Conscientiousness, Extroversion, Agreeableness, and Neuroticism. The models were further applied to the text samples authored by fans of different music genres. Statistical tests reveal significant and reliably detected differences in personality characteristics of individuals with diverse musical preferences. As an example, Classical music enthusiasts show higher levels of agreeableness and lower levels of extroversion, while Hip-Hop fans have lower agreeableness and higher neuroticism. Additionally, Metal fans exhibit a higher propensity for neuroticism, and Electronic music listeners demonstrate greater openness.
Our contribution in this study is, therefore, manyfold: First, we release a carefully collected dataset of authentic linguistic productions of users with diverse and reliably identified musical preferences. Second, we train models for accurate personality detection from text, and release their training data. Finally, through large-scale empirical analysis, we shed a new and interesting light on the association between musical preferences and personality traits, as manifested in our language. All data and code are available at https://github.com/<anonymized_for_submission>.
@@ 2 Related Work
The release of the MyPersonality dataset by (Kosinski et al. [19]—not a non-controversial outcome of data collection through a Facebook application—sparked another line of research, using more advanced methods, inspired by a range of deep-learning architectures [21, 30, 35, 28, 33]. Working with MyPersonality [19], whose support was officially discontinued in 2018, and Essays [25], researches achieved higher accuracy, obtaining around 0.70 on the binary (high or low) personality trait detection task on average.
@@ 3 Dataset
We collected a large and diverse dataset of spontaneous and authentic written content from individuals with distinct musical preferences. We aimed to gather texts that reflect a wide range of commonly discussed topics, deliberately excluding any direct references to music, to avoid potential biases and ensure that the analysis accurately captures the relationship between musical preferences of an individual and their personality traits.
@@ 3.2 Dataset Collection
Given a Reddit UserID of each of the 5,000 authors, the entire user’s digital footprint can be retrieved from the platform, providing direct access to a person’s linguistic productions across a variety of topics, spanning several years. We applied multiple filtering and cleanup steps: (1) excluded subreddits (in)directly related to music [^3] About 150 threads were excluded per manual inspection., (2) aiming at long enough, ideally paragraph-length texts for personality analysis, we filtered out texts shorter than 40 tokens, [^4] The minimal number of tokens is aligned with the mean paragraph length in our training data (see Section 5). and (3) removed duplicate entries. These steps resulted in a clean and high quality dataset; the dataset statistics are reported in Table 1.
@@
Figure 1: Distribution of the number of users (out of the pool) that posted or commented at least once in one of the top-10 most popular subreddits in the data, by genre.
4 Personality Detection from Language:
Datasets and Approaches
@@ 4.1.1 Essays Dataset
Originating from 2,468 essays (collected between 1997 and 2004, each written by a student), these writing submissions were part of a psychology course assignment [25]. Students’ personality scores were determined using the Big Five inventory, through a self-reported questionnaire that measures the five personality traits. Each essay was then labeled with five personality scores, based on a student’s questionnaire, where the mean continuous scores where transformed into binary labels: "yes" (for high-), and "no" (for low-trait presence, respectively).
@@ 4.1.2 MyPersonality Dataset
MyPersonality dataset contains information from a Facebook application originally used by over 6 million users [19]. Similarly to Essays, personality scores were assigned to users based on self-reported personality questionnaire, along with several social network measures, including network size, density, brokerage, and transitivity. The final dataset includes textual Facebook status updates assigned with continuous scores for each personality trait per user. MyPersonality has been particularly useful for research because it combines social media text with personality trait labels, allowing the study of personality in a more naturalistic setting. However, the project was discontinued in 2018: [^5] https://shorturl.at/vYsR9 its data is no longer publicly available, and only a minor subset of 9,917 texts by 250 users is available online.
@@ 4.2 Approaches to Personality Detection
Evaluating performance of the three tools on both datasets yields disappointingly (but not surprisingly) low results. We report a tool’s per-trait accuracy: the ratio of data examples assigned with the correct (binary) label out of the total amount. Considering the random baseline of 0.5, and despite the exposure of the tools to the datasets, the accuracy is low (inline with what is reported in prior art), questioning the applicability of these approaches to the task at hand. Table 2 reports classification accuracy results of existing tools on the two available datasets. We attribute the poor results mainly to the limitations of the existing datasets (see Section 4.1), and anticipate that high quality training data coupled with contemporary modeling, will advance the state-of-the-art in this field. We describe our methodology in the next section.
@@ 5 Methodology
5.1 Big Five Generated Datasets
We applied a systematic approach to generate textual content that could be used to train and evaluate classification models for each of the Big Five personality traits, as described below:
Definition Compilation
@@ 5.1 Big Five Generated Datasets
Prompt Creation
For each personality trait, we wrote a primary prompt that started with the trait’s definition and ended with a request for the LLM to generate several paragraphs demonstrating a high (and similarly, low) level of the trait. As a concrete example, for Extroversion, we generated texts "written by" people who are highly extroverted and then by those who score low on Extroversion, i.e. considered introverts. Appendix A contains our final prompt definitions for the five traits.
We iteratively modified the requests to ensure the generated texts were unique, authentic, and avoided repetition. We also instructed the models to write as if they were different personas, iterating over a variety of topics and writer profiles, e.g., texts written by high school teenagers, by sports lovers (fans or practitioners), politicians, and also texts about finance, news, traveling, etc.
Dataset Generation
Using GPT3.5 for training set collection, and Gemini for test set, we collected hundreds of short passages exhibiting the two extremes of each personality trait. Table 3 reports the final statistics of the collected GenBigFive dataset, after cleanup and filtering, and Table 4 presents a few generated text examples, produced as "high" and "low" trait level by GPT3.5.
@@
Table 4: Example texts generated by GPT3.5 with high and low presence of two personality traits: EXT and AGR.
@@
Human Evaluation
We evaluated the quality of generated data through human annotation of 250 samples: 50 samples were randomly selected for each trait, split equally between low and high. We used the Appen platform [^12] https://www.appen.com/ to recruit native English speakers, who were presented with a personality trait description (identical to what the LLM was prompted with, see Appendix A.2) along with a few examples. Quizzes were integrated into the task to ensure high-quality outcome. Each sample was annotated by five annotators, with an exception of 17 samples, that were annotated only by three. Each example was assigned a trait level (low or high) based on the majority vote. With sufficient training, the task turned out to be relatively easy: annotation agreement with GPT3.5 generation was 94.7% and inter-annotator agreement was kk=0.76.
@@ 5.2 Building New Personality Classifier(s)
We used the collected datasets for each of the five traits for training reliable and accurate personality classifiers. Here, we follow the approach successfully used in prior studies for classification of emotion [1] and the degree of concept abstractness [12], training a logistic regression classifier with passage embeddings as feature vectors. This methodology was chosen for its simplicity, effectiveness, the inherent ability to produce a continuous output score (posterior probability), and efficiency (fast training and inference). We trained five logistic regression models, one per trait, where the passages with low and high trait level were encoded into embeddings using the intfloat/e5-large-v2 encoder [31], [^13] https://huggingface.co/intfloat/e5-large-v2 due to its proven benefits. We trained the models on GPT3.5-generated data (train set) and tested on Gemini-generated data (test set); that, in order to avoid possible confounds in train and test data, generated by the same LLM. We also tested the off-the-shelf tools (see Section 4.2), showing that the newly trained classifiers yield superior results.
@@ 5.2.1 Evaluation Results
We report evaluation accuracy of the five classifiers in Table 5. As a complementary experiment, we also trained the classifiers on the GPT3.5-generated dataset combined with the MyPersonality dataset, and tested on the Gemini-generated test set. High test set accuracy is achieved for each of the five traits, with Openness (OPN) obtaining the lowest accuracy of 0.874. Additional MyPersonality training data harms the results in all cases but CON. Finally, existing tools yielded much inferior, nearly random results.
Table 5: Logistic Regression models’ accuracy, when predicting the Big Five personality traits on the unseen Gemini-generated test set. We report results for two settings: (1) train set comprising solely the GPT3.5-generated data, and (2) + the MyPersonality dataset. Best result in a column is boldfaced.
@@ 6 Experimental Results and Discussion
We next applied the intfloat/e5-large-v2 encoder (see Section 5.2) to over 500K Reddit posts and comments, written by $\sim$5,000 Redditors with five distinct musical preferences, roughly 1,000 for each genre. We used the pre-trained classification models for prediction, assigning each text with five continuous scores, across five personality trait dimensions. Table 6 presents example posts from our dataset, along with their automatically assigned low and high personality scores.
Table 6: Example user texts illustrating varying levels of the Openness (OPN) and Neuroticism (NEU) traits, along with a probability score indicating the likelihood of high trait presence (posterior), as produced by the models.
@@
NEU feel like i’m going crazy. everything is up in the air. no idea where i’ll be in august. i’m applying for jobs and my soul tears apart each time because the process is designed to be fucking excruciating and make you feel worthless (so they can hire you cheaper). […] it is so awful i hate it 0.811
@@
Human Evaluation
We conducted human evaluation, this time for labeling actual Redditors’ content, scored by the classifiers, to make sure the predictive ranking fits human intuition. Similarly to the first task, 250 samples in total where selected for annotation—where the continuous score was binarized into two extremes—this time yielding lower inter-annotator agreement of kk=0.67 and agreement of 86.2% with scores (low or high) assigned by the models. The lower agreement is expected since human annotation is compared to the (imperfect) classifiers outcome.
@@ 6.1 Authors’ Personality Exhibits Association with their Musical Preferences
Next we assigned each user with scores, reflecting their mean personality ranking on each of the Big Five traits. Specifically, we averaged the prediction scores across all texts authored by a given user. Consequently, a community-level personality trait score was computed by averaging over individual scores of users in the specific community. As a concrete example, the mean OPN score of the Classical music lovers was calculated by averaging over the OPN score of the 982 this genre’s fans in our dataset. Figure 2 presents the results. Many of the findings align with common intuition and prior findings in non-linguistic studies [4, 17]: Classical music enthusiasts tend to exhibit the highest levels of Agreeableness, while Hip-Hop fans show the lowest levels, compared to all other genres. Additionally, Classical music enthusiasts generally show lower levels of Extroversion, Hip-Hop and Metal fans tend to have higher levels of Neuroticism.
Figure 2: Mean personality trait by musical genre fans. All five traits exhibit significant within-community differences (pp<0.001), as measured by the ANOVA test.

### v2

# Identity-Aware AI workshop at 28th European Conference on Artificial Intelligence, October 25, 2025, Bologna, Italy

@@ Abstract
Music serves as a powerful reflection of individual identity, often aligning with deeper psychological traits. Prior research has established correlations between musical preferences and personality, while separate studies have demonstrated that personality is detectable through linguistic analysis. Our study bridges these two research domains by investigating whether individuals’ musical preferences leave traces in their spontaneous language through the lens of the Big Five personality traits (Openness, Conscientiousness, Extroversion, Agreeableness, and Neuroticism). Using a carefully curated dataset of over 500,000 text samples from nearly 5,000 authors with reliably identified musical preferences, we build advanced models to assess personality characteristics. Our results reveal significant personality differences across fans of five musical genres. We release resources for future research at the intersection of computational linguistics, music psychology and personality analysis.
keywords:
@@ 1 Introduction
"Our language is the reflection of ourselves. A language is an exact reflection of the character and growth of its speakers."
Cesar Chavez (civil rights activist)
Music is a powerful expression of individuality, often reflecting deeper aspects of one’s character and personality. The relationship between musical preferences and personality has long been a subject of interest in psychology [1, 2, 3, 4]. Prior studies indicate that certain genres are associated with specific personality traits, suggesting that music genre preferences offer valuable insights into individuals’ psychological profile [3, 5, 6].
@@
Our study aims to connect the two lines of research – we explore the relationship between musical preferences and personality traits, as manifested in authentic written language. Specifically, we ask if there are detectable (similarities and) differences in the Big Five personality traits in the language of people with various musical preferences, when authoring spontaneous textual content on social media. We consider five popular and diverse musical genres: Classical music, Hip-Hop, Metal, Indie and Electronic. The hypothesis driving this research posits that despite the inherent complexity of automatic personality analysis, the differences in personality traits of people with various musical preferences "shine through" their language, to the extent that can be automatically captured through advanced natural language processing techniques.
We explore the hypothesis on a carefully curated large and diverse dataset, consisting of over 500,000 text samples written by nearly 5,000 authors with reliably identified musical preferences. These samples were collected from a range of public non-music-related online forums, ensuring a broad topical representation. Each participant was linked to a single dominant music genre they frequently engage with, enabling unbiased textual personality extraction and a focused analysis of the relationship between personality traits and musical preferences. Leveraging text generation capabilities of contemporary LLMs we then gather high-quality data for training personality detectors along the five dimensions: Openness, Conscientiousness, Extroversion, Agreeableness, and Neuroticism.
The models were further applied to the text samples authored by fans of different music genres. Statistical tests reveal significant and reliably detected differences in personality characteristics of individuals with diverse musical preferences. As an example, Classical music enthusiasts show higher levels of Agreeableness and lower levels of Extroversion, while Hip-Hop fans have lower Agreeableness and higher Neuroticism. Additionally, Metal fans exhibit a higher propensity for Neuroticism, and Electronic music listeners show greater Openness.
Our contribution in this study is, therefore, manyfold: First, we release a carefully collected dataset of authentic linguistic productions of users with diverse and reliably identified musical preferences. Second, we train models for accurate personality detection from text, and release their training data. Finally, through large-scale empirical analysis, we shed a new and interesting light on the association between musical preferences and personality traits, as manifested in our language. All our data and code are available at https://github.com/eliranshemtov/Musical-Preferences-And-Textual-Expression.
@@ 2 Related Work
The release of the MyPersonality dataset by Kosinski et al. [17] — not a non-controversial outcome of data collection through a Facebook application — sparked another line of research, using more advanced methods, inspired by a range of deep-learning architectures [18, 9, 19, 20, 21]. Working with MyPersonality [17], whose support was officially discontinued in 2018, and Essays [15], researches achieved higher accuracy: around 0.70 on the binary (high or low) personality trait detection task.
@@ 3 Dataset
We collected a large and diverse dataset of spontaneous and authentic written content from individuals with distinct musical preferences. We aimed to gather texts that reflect a wide range of commonly discussed topics, deliberately excluding any direct references to music, to avoid potential confounds and ensure that the analysis accurately captures the relationship between musical preferences of an individual and their personality traits.
@@ 3.2 Dataset Collection
Given a Reddit UserID of each of the 5,000 authors, the entire user’s digital footprint can be retrieved from the platform, providing direct access to a person’s linguistic productions across a variety of topics, spanning several years. We applied multiple filtering and cleanup steps: (1) excluded subreddits (in)directly related to music, [^3] About 150 threads were excluded per manual inspection. (2) aiming at long enough, ideally paragraph-length texts for personality analysis, we filtered out texts shorter than 40 tokens, [^4] The minimal number of tokens is aligned with the mean paragraph length in our training data (see Section 5). and (3) removed duplicate entries. These steps resulted in a clean and high quality dataset; the dataset statistics are reported in Table 1.
@@
Figure 1: Distribution of the number of users (out of the pool) that posted or commented at least once in one of the top-10 most popular subreddits in the data, by genre. The various topical threads are represented roughly equally by the fans of the five musical genres.
4 Personality Detection from Language: Datasets and Approaches
@@
Importantly, we conclude that the existing annotated training datasets for personality classifier are not sufficiently adequate for the task, suffering from various drawbacks, as detailed below.
@@ 4.1.1 Essays Dataset
Originating from 2,468 essays (collected between 1997 and 2004, each written by an undergraduate student), these writing submissions were part of a psychology course assignment [15]. Students’ personality scores were determined using the Big Five inventory, through a self-reported questionnaire that measures the five personality traits. Each essay was then labeled with five personality scores, based on a student’s questionnaire, where the mean continuous scores where transformed into binary labels: "yes" (for high-), and "no" (for low-trait presence, respectively).
@@ 4.1.2 MyPersonality Dataset
MyPersonality dataset contains information from a Facebook application originally used by over 6 million users [17]. Similarly to Essays, personality scores were assigned to authors based on self-reported personality questionnaire, along with several social network measures, including network size, density, brokerage, and transitivity. The final dataset includes textual Facebook status updates assigned with continuous scores for each personality trait per author. MyPersonality has been particularly useful for research because it combines social media text with personality trait labels, allowing the study of personality in a more naturalistic setting. However, the project was discontinued in 2018 due to the heavy burden of its maintenance: [^5] https://sites.google.com/michalkosinski.com/mypersonality its data is no longer publicly available, and only a minor subset of 9,917 texts by 250 users is available online.
@@ 4.2 Approaches to Personality Detection
Evaluating performance of the three tools on both datasets yields disappointingly (but not surprisingly) low results. We report a tool’s per-trait accuracy: the ratio of data examples assigned with the correct (binary) label out of the total amount. Considering the random baseline of 0.5, and despite the exposure of the tools to the datasets, the accuracy is low (inline with what is reported in prior art), questioning the applicability of these approaches to the task at hand.
Table 2 reports classification accuracy results of existing tools on the two available datasets. We attribute the poor results mainly to the limitations of the existing datasets (see Section 4.1), and anticipate that high quality training data coupled with contemporary modeling, will advance the state-of-the-art in this field. We describe our approach and methodology in the next section.
@@ 5 Methodology
5.1 Big Five Generated Datasets (GenBigFive)
We applied a systematic approach to generate textual content that could be used to train and evaluate classification models for each of the Big Five personality traits, as described below.
5.1.1 Definition Compilation
@@
5.1.2 Prompt Creation
For each personality trait, we wrote a primary prompt that started with the trait’s definition and ended with a request for the LLM to generate several paragraphs demonstrating a high (and similarly, low) level of the trait. As a concrete example, for Extroversion, we generated texts "written by" people who are highly extroverted and then by those who score low on Extroversion, i.e. considered introverts. Appendix A contains our final prompt definitions for the five traits. We iteratively modified the requests to ensure the generated texts were unique, authentic, and avoided repetition. We also instructed the models to write as if they were different personas, iterating over a variety of topics and writer profiles, e.g., texts written by high school teenagers, by sports lovers (fans or practitioners), politicians, and also texts about finance, news, traveling, etc.
5.1.3 Dataset Generation
Using GPT3.5 for training set collection, and Gemini for test set, we collected hundreds of short passages exhibiting the two extremes of each personality trait. Table 3 reports the final statistics of the collected GenBigFive dataset, after cleanup and filtering, and Table 4 presents a few generated text examples, produced as exhibiting "high" and "low" trait level by GPT3.5.
@@ 5.1.3 Dataset Generation
Table 4: Example texts generated by GPT3.5 with high and low presence of two personality traits: EXT and AGR. The generated texts are purposefully exaggerated (by crafting prompts accordingly) towards high or low trait requirements, for training classifiers, which benefit from extreme examples.
@@
5.1.4 Human Evaluation
We evaluated the quality of generated data through human annotation of 250 samples: 50 samples were randomly selected for each trait, split equally between low and high. We used the Appen platform [^12] https://www.appen.com/ to recruit native English speakers, who were presented with a personality trait description (identical to what the LLM was prompted with, see Appendix A.2) along with a few examples. Quizzes were integrated into the task to ensure high-quality outcome. Each sample was annotated by five annotators, with an exception of 17 samples, that were annotated only by three. Each example was assigned a trait level (low or high) based on the majority vote.
With sufficient training, the task turned out to be relatively easy: annotation agreement with GPT3.5 generation was 94.7% and inter-annotator agreement was kk=0.76.
@@ 5.2 Building New Personality Classifier(s)
We used the collected datasets for each of the five traits for training reliable and accurate personality classifiers. Here, we follow the approach successfully used in prior studies for classification of emotion [29] and the degree of concept abstractness [30], training a logistic regression classifier with passage embeddings as feature vectors. This methodology was chosen for its simplicity, effectiveness, the inherent ability to produce a continuous output score (posterior probability), and efficiency (fast training and inference).
We trained five logistic regression models, one per trait, where the passages with low and high trait level were encoded into embeddings using the intfloat/e5-large-v2 encoder [31], [^13] https://huggingface.co/intfloat/e5-large-v2 due to its proven benefits on multiple tasks. We trained the models on GPT3.5-generated data (train set) and tested on Gemini-generated data (test set); that, in order to avoid possible confounds in train and test data, generated by the same LLM. We also tested the off-the-shelf tools (see Section 4.2), showing that the newly trained classifiers yield superior results on the unseen test set.
@@ 5.2.1 Evaluation Results
We report evaluation accuracy of the five classifiers in Table 5. As a complementary experiment, we also trained the classifiers on the GPT3.5-generated dataset combined with the MyPersonality dataset, and tested on the Gemini-generated test set. High test set accuracy is achieved for each of the five traits, with Openness (OPN) obtaining the lowest accuracy of 0.874. Additional MyPersonality training data harms the results in all cases but CON. Finally, existing tools yielded inferior, nearly random results.
Table 5: Our classifiers’ accuracy, when predicting the Big Five personality traits on the unseen Gemini-generated test set. We report results for two settings: (1) train set comprising solely the GPT3.5-generated data, and (2) + the MyPersonality dataset. Best result in a column is boldfaced.
@@ 6 Experimental Results and Discussion
We next applied the intfloat/e5-large-v2 encoder (see Section 5.2) to over 500K Reddit posts and comments, written by nearly 5,000 Redditors with five distinct musical preferences, roughly 1,000 for each genre. We used the pre-trained classification models for prediction, assigning each text with five continuous scores, across five personality trait dimensions.
Table 6 presents example posts from our dataset, along with their automatically assigned low and high personality scores.
Table 6: Example user texts (verbatim from the dataset) illustrating varying levels of the Openness (OPN) and Neuroticism (NEU) traits, along with a probability score indicating the likelihood of high trait presence (posterior), as produced by the respective regression models.
@@
NEU feel like i’m going crazy. everything is up in the air. no idea where i’ll be in august. i’m applying for jobs and my soul tears apart each time because the process is designed to be fucking excruciating and make you feel worthless (so they can hire you cheaper). […] 0.811
@@
6.1 Human Evaluation
We conducted human evaluation, this time for labeling actual Redditors’ content, scored by the classifiers, to make sure the predictive ranking fits human intuition. Similarly to the first task, 250 samples in total where selected for annotation — where the continuous score was binarized into two extremes — this time yielding lower inter-annotator agreement of kk=0.67 and agreement of 86.2% with scores (low or high) assigned by the models. We note that the lower agreement (recall the 94.7% for generated content) is expected since human annotation is compared to the (imperfect) classifiers outcome.
@@ 6.2 Authors’ Personality Exhibits Association with their Musical Preferences
Next we assigned each user with scores, reflecting their mean personality ranking on each of the Big Five traits. Specifically, we averaged the prediction scores across all texts authored by a given user. Consequently, a community-level personality trait score was computed by averaging over individual scores of users in the specific community. As a concrete example, the mean OPN score of the Classical music lovers was calculated by averaging over the OPN score of the 982 this genre’s fans in our dataset.
Figure 2 presents the results. Many of the findings align with common intuition and prior results in non-linguistic studies [6, 12]: Classical music enthusiasts tend to exhibit the highest levels of Agreeableness, while Hip-Hop fans show the lowest levels, compared to all other genres. Additionally, Classical music enthusiasts generally show lower levels of Extroversion, Hip-Hop and Metal fans tend to have higher levels of Neuroticism.
Figure 2: Mean personality trait by musical genre fans. All five traits exhibit significant within-community differences, as measured by the ANOVA test.